OpenStudy (anonymous):

Not sure where to start: Solve for Tf [1.0][100-Tf][0.50] = [50.0][Tf-20][4.184 ]

OpenStudy (anonymous):

Not sure where to start: \[ Solve for T _{f} [1.0][100-T _{f}][0.50] = [50.0][T _{f}-20][4.184 ]\]

OpenStudy (anonymous):

what do you meant by \[T_f\] here

OpenStudy (anonymous):

Well, I typed it weird. All the info I have is to solve for Tf (with the f being a sub). So the real problem is \[[1.0][100−Tf][0.50]=[50.0][Tf−20][4.184]\]

OpenStudy (anonymous):

what is T and F here then?

OpenStudy (anonymous):

\[[1.0][100−T _{f}][0.50]=[50.0][T _{f}−20][4.184]\]

OpenStudy (phi):

Tf probably stands for Temperature_final are the square braces really just parentheses?

OpenStudy (phi):

if we use x for Tf (100-x) * ½ = 50*4.184 *(x-20) 100-x = 100*4.184(x-20) 100 - x = 481.4x - 20*418.4 add x to both sides 100 = 482.4x - 20*418.4 add 20*418.4 to both sides 100+20*418.4= 482.4 x divide both sides by 482.4

OpenStudy (anonymous):

It doesn't say. Simply says solve for Tf. I think it's just an exercise ... I'm not sure about the square braces. They are there as square braces. Don't know if she means ( ).

OpenStudy (anonymous):

are you studying algebra?or something I think It's just asking to solve for Tf

OpenStudy (anonymous):

Yes, algebra

OpenStudy (anonymous):

DO you expect this ..: Then here is the answer \[\frac{ 100 - T_f }{ 2 } = [50T_f - 1000][4.184]\] \[100-T_f = 2([50T_f -1000] [4.184])\] carrryibng LHS?RHs working we can work out T_f

OpenStudy (anonymous):

Not sure where you got the 1000, Adjax??

OpenStudy (anonymous):

The water and the metal will end at the same temperature, Tf (T final) The heat lost by the metal will be equal to the heat gained by the water. The heat gained(or lost) = specific heat x mass x temperature change Heat lost by metal = 0.50 J/gC x 1000 g x (100 - Tf) (0.50 kJ/kgC = 0.50 J/gC) Heat gained by water = 4.184 J/gC x 50.0 x (Tf - 20.0) 0.50 x 1000 x (100 - Tf) = 4.184 x 50.0 x (Tf - 20) 50000 - 500Tf = 209.2Tf - 4184 54184 = 709.2 Tf Tf = 76.4 C

OpenStudy (anonymous):

did that help you

OpenStudy (anonymous):

Don't see how you got to the 2nd line from the 1st line...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

A 1.0 kg sample of metal with a specific heat of 0.50 J/g (degrees) C is heated to 100 (degrees) C and then placed in a 50.0 g sample of water at 20.0 (degrees) C. What is the final temperature of the metal and the water?

OpenStudy (anonymous):

that shold help you if not chat me

OpenStudy (anonymous):

@gra first toy tell us that you are studying algebra/maths ..now here's a solution ABOVE OF PHYSICS WHICH IS ALSO RIght in it's frame of reference.. So for what you are asking solution for?

OpenStudy (anonymous):

@cece1902 .. not sure why this is so difficult for me to understand! I think the teacher is simply wanting us to go through the algebraic steps...but I'm having trouble setting it up.

OpenStudy (anonymous):

ok just put this.A 1.0 kg sample of metal with a specific heat of 0.50 J/g (degrees) C is heated to 100 (degrees) C and then placed in a 50.0 g sample of water at 20.0 (degrees) C. What is the final temperature of the metal and the water?

OpenStudy (anonymous):

did that help you

OpenStudy (anonymous):

or not